OpenStudy (anonymous):

√((y^15z^3)/(9))(√((81x^8)/(x^6y^9z^3)))

OpenStudy (anonymous):

\[\sqrt{\frac{ y^15z^3 }{ 9 }}\ \times \sqrt{\frac{ 81x^8 }{ x^6y^9z^3? }}\] this is the question

OpenStudy (anonymous):

oh my god,difficult to type

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):